It costs m dollars to buy 48 strawberries, at the same rate, how many dollars will it cost to buy 30 strawberries?
Answer by stanbon(75887)

It costs m dollars to buy 48 strawberries, at the same rate, how many dollars will it cost to buy 30 strawberries?
----
Use a proportion::
x/30 = m/48
x = 30m/48 = (5/8)m dollars
-------------
Cheers,
Stan H.
---------